Circular No. 5846

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993p) Jean Mueller reports her discovery of a comet on a plate taken by

J. D. Mendenhall and herself with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. The object has a condensed nuclear region and a tail fanned to the west-northwest. A confirming film was taken by E. and C. Shoemaker and D. Levy with the 0.46-m Palomar Schmidt telescope on Aug. 17. The following measurements are by Mueller:

1993 UT α_{2000} δ_{2000} m_{1} Observer
Aug. 16.44375 1^{h}45^{m}29.^{s}07 +49^{o}49′08′′.1 14.5 Mueller
16.49444 1 45 27.92 +49 49 15.3  _{”}
17.49253 1 44 56.20 +49 51 42.0 14  Shoemaker

NOVA AQUILAE 1993 C. E. Woodward and J. Cole, University of Wyoming, report the fol-

lowing infrared broadband magnitudes, obtained using the WIRO 2.3-m telescope (+ University of Minnesota bolometer) on Aug. 1.3 UT, H = 8.19 ± 0.14, K = 5.81 ± 0.12, L = 2.89 ± 0.10, M = 1.95 ± 0.33; Aug. 7.2, J = 9.84 ± 0.07, H = 8.83 ± 0.04, K = 5.84 ± 0.01, L = 3.05 ± 0.02, M = 1.50 ± 0.01, N = 0.05 ± 0.01.

Visual magnitude estimates by W. G. Dillon, Missouri City, TX: July 24.27 UT, 14.0; Aug. 12.34, 13.6.

V992 SCORPII P. Camilleri, Cobram, Victoria, Australia, reports additional predis-

covery photovisual magnitude data of Nova Sco 1992 (cf. IAUC 5526, 5551, etc.): 1992 Apr. 30.65 UT, [12 (Camilleri); May 4.51, 8.2 (R. T. Price, Bethanga, Victoria).

SUPERNOVA 1993J IN NGC 3031 Further photoelectric photometry by H. Mikuž, B. Dintinjana, and T.

Zwitter, Ljubljana, Slovenia (cf. IAUC 5796): July 23.86 UT, B = 14.25 ± 0.02, V = 13.62 ± 0.01, R = 13.19 ± 0.01, I = 12.76 ± 0.02; Aug. 14.08, 14.73 ± 0.05, 14.08 ± 0.02, 13.79 ± 0.02, 13.31 ± 0.02.

1993 August 18 Daniel W. E. Green